OpenStudy (anonymous):

The time required to complete a task VARIES INVERSELY with the number of people working on the task. It takes 6 people 40 hours to complete a task. How long will it take 20 people to complete the same task?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

My guess is 3 hours?

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

@pooja195

OpenStudy (igreen):

Hold on

OpenStudy (igreen):

No, I think that's wrong..

OpenStudy (igreen):

Set up a proportion. \(\dfrac{6}{40} = \dfrac{20}{x}\)

OpenStudy (igreen):

Cross multiply and solve for 'x'

OpenStudy (anonymous):

6 times X 20 times 40

OpenStudy (anonymous):

so you have 6x and 800

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

Yep, 6x = 800 Divide 6 to both sides, what's 800 / 6?

OpenStudy (anonymous):

0.0075

OpenStudy (igreen):

No, check again.

OpenStudy (anonymous):

133?

OpenStudy (anonymous):

@igreen

OpenStudy (igreen):

Oh wait, nevermind..yes..

OpenStudy (anonymous):

???

OpenStudy (igreen):

Okay, I did something wrong..hold on.

OpenStudy (anonymous):

So 133 minutes?

OpenStudy (igreen):

No

OpenStudy (anonymous):

erm :?

OpenStudy (igreen):

Formula for inverse variation is: \(y = \dfrac{k}{x}\) Plug in 6 for 'x' and '40' for 'y'. \(40 = \dfrac{k}{6}\)

OpenStudy (igreen):

Multiply 6 to both sides, what's 40 * 6?